OpenStudy (anonymous):

if x is positive and not equal to one, simplify the sqrt of x divided by x^3

OpenStudy (jmark):

Given if x is positive and not equal to one sqrt(x)/x^3 = (x^1/2)/x^3 = x^(1/2-3) =x^(-5/2) =1/x^5/2 Know more about Completing the Square Calculator http://goo.gl/QBy04r